OpenStudy (anonymous):

Does there exist an infinite sequence \(p_0,p_1,p_2,... \) of prime numbers such that \(p_k\in\{4p_{k-1}-1,\,4p_{k-1}+1\}\qquad (p_k=4p_{k-1}\pm1)\) for all \(k\in\mathbb{Z}^+ \)

ganeshie8 (ganeshie8):

interesting, im thinking if its related to http://en.wikipedia.org/wiki/Dirichlet's_theorem_on_arithmetic_progressions

ganeshie8 (ganeshie8):

may be not as the next term could be 4p+1 or 4p-1 so that kills the arithmetic progression

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

i have a feeling that there does not exist , but still trying to prove

ganeshie8 (ganeshie8):

lets see whats going on, start with \(\large p_0 = 3\) : \(\large 4*3\pm 1 \) both are primes, pick \(\large p_1 = 11\) \(\large 4*11- 1 \) is a prime, pick \(\large p_2 = 43\) \(\large 4*43+1 \) is a prime, pick \(\large p_3 = 173\)

OpenStudy (anonymous):

brb

myininaya (myininaya):

This might be hard to prove or find. Isn't it just a conjecture that there are infinitely many primes of the form 4p+1 .

ganeshie8 (ganeshie8):

Clearly it is flip flopping - it has to alternate because, if two consecutive terms are of form 4p+1, then one of them is divisible by 3 : \[\large 4(4p+1)+1 \equiv p+2 \pmod 3 \] if \(p \equiv 1 \pmod 3\) then p+2 is divisible by 3 - contradiction if \(p \equiv -1 \pmod 3\) then ...

ganeshie8 (ganeshie8):

not sure 100% of above conclusion, need to tweak a bit..

ganeshie8 (ganeshie8):

the previous sequence ends at \(p_5 = 2765\) http://www.wolframalpha.com/input/?i=is+%284*691%2B1%29+a+prime

ganeshie8 (ganeshie8):

so we get a recurrence relation : \[\large a_{n+1} = 4a_n + (-1)^n\] solving it gives : \[\large a_n = a_04^{n-1} + \dfrac{(-1)^n-4^n}{5}\]

ganeshie8 (ganeshie8):

the problem simplifies to proving one of these terms is composite

ganeshie8 (ganeshie8):

you can conlude like this i hope : \(\large 4^{n-1} \equiv \pm 1 \pmod 5\) so for some \(n\), the expression \(\large \dfrac{(-1)^n-4^n}{5}\) becomes \(\large \mp 1\mod 5 \) making that term divisible by 5 \(\huge \square \)

OpenStudy (anonymous):

Brilliant !

OpenStudy (anonymous):

@myininaya 4p+3 or 4p+1 yes its true it gives infinitly many primes , but im talking about a sequense of p's such that blah blah blah

OpenStudy (ikram002p):

*

ganeshie8 (ganeshie8):

looking at about above proof today again freshly - I see two silly mistakes/blunders

ganeshie8 (ganeshie8):

challenge question : figure out the two mistakes :)

OpenStudy (ikram002p):

haha xD

ganeshie8 (ganeshie8):

they can be corrected easily thouhg. .

OpenStudy (ikram002p):

cant see :'(

ganeshie8 (ganeshie8):

guess i should tell about the 2nd mistake : \[\large \begin{align}a_n &= a_04^{n-1} + \dfrac{(-1)^n-4^n}{5} \\&\equiv \pm a_0 + \dfrac{(-1)^n-4^n}{5} \pmod 5 \end{align} \]

OpenStudy (ikram002p):

give up xD ugh tell

ganeshie8 (ganeshie8):

that first prime, \(\large a_0\) sticks in there.

ganeshie8 (ganeshie8):

so we need to show that above expression gives all the residues in `mod 5` to conclude the proof

ganeshie8 (ganeshie8):

simply saying +-1 and -+1 wont do. need to prove it..

OpenStudy (ikram002p):

ok then xD

ganeshie8 (ganeshie8):

is there an easy way to prove that the expression \(\large\dfrac{(-1)^n-4^n}{5}\) leaves all the residues in mod 5 ?

ganeshie8 (ganeshie8):

actually we can forget about \(\large a_0\) because its just a constant..

ganeshie8 (ganeshie8):

if the set of elements {x} leave all the residues in mod 5, {x+c} also leaves all the residues

ganeshie8 (ganeshie8):

so proving \(\dfrac{(-1)^n-4^n}{5}\) leaves all the residues in mod 5 is sufficient i think...

OpenStudy (ikram002p):

hmm im not sure :O

ganeshie8 (ganeshie8):

maybe we can work n=0,1,2,.. and see what they evaluate to

ganeshie8 (ganeshie8):

\[\large \dfrac{(-1)^1-4^1}{5} \equiv 4\pmod 5\] \[\large \dfrac{(-1)^2-4^2}{5} \equiv 2\pmod 5\] \[\large \dfrac{(-1)^6-4^6}{5} \equiv 1\pmod 5\] etc... im not sure how to prove it without listen them all down at the moment..

OpenStudy (ikram002p):

hmm well i guess its enough :O

ganeshie8 (ganeshie8):

ugh ok il try again when i get free time

OpenStudy (ikram002p):

mmm ok